# Google Research

### v1

@@ Abstract
[…]
@@ 1 Introduction
Recently, Tamkin et al. (2024) proposed Clio, a system that uses LLMs to aggregate insights while preserving user privacy; building on Zheng et al. (2023); Zhao et al. (2024) (see §D for more details). Clio provides insights about how LLM chatbots (Claude.ai in their case) are used in the real world and visualizes these patterns in a graphical interface. Subsequently, Handa et al. (2025) applied Clio to more than four million Claude.ai conversations, to provide insight into which economic tasks are performed by or with chatbot help.
@@
with formal end-to-end differential privacy guarantees?
Differential privacy (DP) (Dwork et al., 2006) is a powerful mathematical privacy notion that is considered the gold standard for privacy protection in data analytics and machine learning, and had been widely applied in practice  (see, e.g., Desfontaines, 2021).
[…]
@@
[…]
@@
[…]
  • In §E:
[…]
@@ 2 Formal Problem and Preliminaries
The input is a dataset of conversations $\bm{x}=\{x_{1},\ldots,x_{n}\}$, where each element $x_{i}$ represents a conversation between a user and an LLM.
@@
The desired output for our problem is $\bm{s}=\{s_{1},s_{2},\ldots,s_{m}\}\in\mathcal{S}^{*}$ consisting of summaries $s_{j}\in\mathcal{S}$. The set $\mathcal{S}$ is potentially infinite (e.g., the set of all strings in our case). The DP guarantee applies only to this output set s of summaries.
@@
Evaluating the utility of privacy-preserving summaries presents unique challenges due to the unstructured nature of text data. For utility measurement purposes only, we produce a mapping $\varphi:\bm{x}\to\bm{s}$ so that $\varphi(x_{i})$ is the summary with which the record $x_{i}$ is associated. Importantly, this mapping violates DP and is never returned—it serves solely as an internal mechanism for evaluation.
@@ 2.1 Differential Privacy
All the datasets in this paper are collections of records. A randomized function that maps input datasets to an output space is referred to as a mechanism. For mechanism $\mathcal{M}$ and dataset $\bm{x}$, we use $\mathcal{M}(\bm{x})$ to denote the random variable returned over the output space of $\mathcal{M}$. Two datasets $\bm{x}$ and $\bm{x}^{\prime}$ are said to be adjacent, denoted $\bm{x}\sim\bm{x}^{\prime}$, if, one can be obtained by adding or removing one record from the other. We consider the following notion of $(\varepsilon,\delta)$-differential privacy (DP).
@@
Let $\varepsilon>0$ and $\delta\in[0,1]$. A mechanism $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-differentially privacy ($(\varepsilon,\delta)$-DP for short) if for all adjacent datasets $\bm{x}\sim\bm{x}^{\prime}$, and for any (measurable) event $E$ it holds that $\Pr[\mathcal{M}(\bm{x})\in E]\leq e^{\varepsilon}\Pr[\mathcal{M}(\bm{x}^{\prime})\in E]+\delta.$
@@
Clustering is a central primitive in unsupervised machine learning (see, e.g., Ikotun et al., 2023).Abstractly speaking, a (Euclidean) clustering algorithm takes as input a dataset of vectors in $\mathbb{R}^{d}$, and returns a set $C\subseteq\mathbb{R}^{d}$ of $k$ “cluster centers” (for a specified parameter $k$), that are representative of the dataset; formalized for example, via the $k$-means objective (Lloyd, 1982). Clustering algorithms have been well-studied in the literature for several decades, and more recently with DP guarantees as well. We discuss more related work in §D. While our framework is compatible with any DP clustering algorithm, specifically, we use the implementation of DP clustering in the open-source Google DP library [^2] https://github.com/google/differential-privacy/tree/main/learning/clustering , as described in Chang & Kamath (2021); we denote the $(\varepsilon,\delta)$-DP instantiation of any such algorithm producing (up to) $k$ clusters as $\texttt{DP\text{-}K\text{-}Means}_{k,\varepsilon,\delta}$.
@@
In the private histogram release problem over a set of bins $\mathcal{B}$, the input is a multi-set $\bm{h}:=\{h_{1},h_{2},\ldots,h_{k}\}$ where each $h_{i}\subseteq\mathcal{B}$ with $|h_{i}|\leq k$. The desired output is a histogram $h_{*}\in\mathbb{Z}^{\mathcal{B}}$, that is desired to be as close to $\sum_{i}h_{i}$ as possible, where we abuse notation to interpret $h_{i}$ as an indicator vector in $\{0,1\}^{\mathcal{B}}$.
@@
In the private partition selection problem over a (potentially infinite) set $\mathcal{B}$, the input is a multi-set $\bm{h}:=\{h_{1},h_{2},\ldots,h_{k}\}$ where each $h_{i}\subseteq\mathcal{B}$ with $|h_{i}|\leq k$. The desired output is a subset $S\subseteq\mathcal{B}$ such that $S\subseteq\bigcup_{i}h_{i}$ and $S$ is as large as possible. We use the partition selection algorithm introduced by Desfontaines et al. (2022); $\texttt{PPS}_{k,\varepsilon,\delta}$ is the instantiation of this algorithm that is $(\varepsilon,\delta)$-DP. See §A for formal details.
@@ 3 Simple-Clio
We present a non-DP (public) framework for hierarchical text summarization, which is a simplified version of Clio (Tamkin et al., 2024). Our approach transforms a dataset $\bm{x}=(x_{1},x_{2},\ldots,x_{n})\in\mathcal{X}^{n}$ of conversations into high-level summaries $\bm{s}=(s_{1},\ldots,s_{m})\in\mathcal{S}^{m}$ through three stages. The pseudocode can be found in Algorithm 4 with the implementation details as discussed below.
@@
    [Embedding Generation] To use a clustering algorithm, we convert conversations to numerical vectors. A typical way for doing so is to use pre-trained embedding models, specifically we use all-mpnet-base-v2 (Reimers & Gurevych, 2019; 2022). However, following Tamkin et al. (2024), we first map the conversation to a summary text before applying the embedding model; this ensures that conversations that are similar in content but structured differently get mapped to embeddings that are closer. We use an LLM to map each conversation to the summary text, using a prompt as described in §F.1.
@@
    [Clustering] We apply the standard $k$-means clustering (Lloyd, 1982) to group similar conversations based on their embeddings. We select $k$ to achieve an average cluster size of approximately 150 conversations. We assume $\textsc{KMeans}_{k}(\bm{e})$ returns the centers of cluster and its assignments (points in that cluster).
@@
    [Summary Generation] For each cluster $\bm{C}_{j}$, we generate a representative summary by sampling both random conversations from the cluster and contrastive conversations near the cluster centroid. We then use an LLM to produce a summary that captures the central theme of the cluster. The prompt used for this step is provided in §F.2.
[…]
@@
Input: Input dataset $\bm{x}=(x_{1},x_{2},\ldots,x_{n})\in\mathcal{X}^{n}$, keyword set $\mathcal{K}$
@@
$\bm{e}\leftarrow\{\mathsf{ExtractEmbedings}(x_{i}):i\in[n]\}$
@@
             $\bm{K}_{j}\leftarrow\text{RandomKeywords}(\mathcal{K},t)$ ;
@@
             $\mathbf{samples}_{j}\leftarrow\text{RandomSample}(\bm{C}_{j},m)$ ;
@@
[…]
@@
             $\bm{K}_{j}\leftarrow\text{TopKeywords}(\mathbf{private\_hist}_{j},t)$ ;
@@
$\bm{s}\leftarrow\{\mathsf{LLMSummarize}(\bm{K}_{j}):j\in[k]\}$
@@
[…]
Our approach takes a dataset $\bm{x}=(x_{1},x_{2},\ldots,x_{n})\in\mathcal{X}^{n}$ of conversations, a predefined set $\mathcal{K}$ of keywords (discussed in §4.1), a number $k\in\mathbb{N}$ of clusters, a number of keywords used for summary, and privacy parameters $\varepsilon_{c},\varepsilon_{\mathrm{hist}},\varepsilon_{\mathrm{size}}>0$ and $\delta_{c}\in[0,1]$. It produces high-level topical summaries $\bm{s}=(s_{1},\ldots,s_{m})\in\mathcal{S}^{m}$ through the following privacy-preserving pipeline:
@@
    [Embedding Generation] This step is the same as in the public pipeline.
@@
    [Clustering] Non-private clustering algorithms usually output both cluster centers and cluster assignments. However, cluster assignments are non-private in the add-remove model; hence, DP clustering algorithms output only cluster centers.
@@
      First, we generate cluster centers using an $(\varepsilon_{\mathrm{c}},\delta_{\mathrm{c}})$-DP k-means algorithm. Although we use the algorithm from Chang & Kamath (2021), any DP clustering algorithm can be used here and the same privacy guarantees would follow.
@@
    • Apply DP to the cluster size with privacy parameter $\varepsilon_{\text{size}}$. If the noisy size falls below a threshold $\tau$, we skip the cluster to prevent privacy leakage from small groups.
@@
    • Use an LLM to extract relevant keywords from the predefined set $\mathcal{K}$ for each sampled conversation. For the purpose of DP, we restrict the keywords chosen from each conversation, say 5 keywords for each conversation. See Prompts in Section F.3.
@@
    [Summary Generation] For each cluster, we generate a high-level summary using only the selected keywords. See Prompts in Section F.4
@@
[…]
@@
The first part satisfies $(\varepsilon_{\mathrm{c}},\delta_{\mathrm{c}})$-DP since it consists of operations applied per record and after that application of the $(\varepsilon_{\mathrm{c}},\delta_{\mathrm{c}})$-DP instantiation of the DP-KMeans algorithm. The second part is $\varepsilon_{\mathrm{size}}$-DP provided that the set of cluster centers is fixed since it again applies per-record operations and an $\varepsilon_{\mathrm{size}}$-DP histogram release algorithm. The third part is $\varepsilon_{\mathrm{hist}}$-DP provided that the set of cluster centers is fixed since it again applies per-record operations and an $\varepsilon_{\mathrm{hist}}$-DP histogram release algorithm, by parallel composition. Altogether, this means that the algorithm producing $\bm{K}_{1}$, …, $\bm{K}_{k}$ is $(\varepsilon_{\mathrm{c}}+\varepsilon_{\mathrm{hist}}+\varepsilon_{\text{size}},\delta_{\mathrm{c}})$-DP due to basic composition.

### v2

# Urania: Differentially Private Insights into AI Use

@@ Abstract
We introduce Urania, a novel framework for generating insights about LLM chatbot interactions with rigorous differential privacy (DP) guarantees. The framework employs a private clustering mechanism and innovative keyword extraction methods, including frequency-based, TF-IDF-based, and LLM-guided approaches. By leveraging DP tools such as clustering, partition selection, and histogram-based summarization, Urania provides end-to-end privacy protection. Our evaluation assesses lexical and semantic content preservation, pair similarity, and LLM-based metrics, benchmarking against a non-private method inspired by Clio (Tamkin et al., 2024). Moreover, we develop a simple empirical privacy evaluation that demonstrates the enhanced robustness of our DP pipeline. The results show the framework’s ability to extract meaningful conversational insights while maintaining stringent user privacy, effectively balancing data utility with privacy preservation.
@@ 1 Introduction
Recently, Tamkin et al. (2024) proposed Clio, a system that uses LLMs to aggregate insights while preserving user privacy, building on Zheng et al. (2023); Zhao et al. (2024) (see §E for more details). Clio provides insights about how LLM chatbots (Claude.ai in their case) are used in the real world and visualizes these patterns in a graphical interface. Subsequently, Handa et al. (2025) applied Clio to more than four million Claude.ai conversations, to provide insight into which economic tasks are performed by or with chatbot help.
@@
with formal end-to-end DP guarantees?
Differential privacy (DP) (Dwork et al., 2006) is a powerful mathematical notion that is considered the gold standard for privacy protection in data analytics and machine learning, and had been widely applied in practice  (see, e.g., Desfontaines, 2021).
Figure 1: Illustration of our proposed Urania method.
@@
We propose Urania, a novel framework for summarizing user queries to an LLM chatbot with a formal end-to-end DP guarantee, that does not rely on any heuristic properties of the underlying LLM used (§4); our approach uses LLMs for keyword extraction, and utilizes existing tools including DP clustering, partition selection and histogram release (see §2 for descriptions of these tools). En route,
@@
    we propose a family of evaluation metrics to compare the quality of any two query summarization methods, and we use these to compare the performance of Simple-Clio and Urania, and finally,
  • In §F:
    we propose a simple empirical privacy evaluation method using a membership inference-style attack to conceptually demonstrate that indeed Urania is more robust than Simple-Clio.
@@ 2 Formal Problem and Preliminaries
The input is a dataset $\bm{x}=\{x_{1},\ldots,x_{n}\}$ of conversations, where each element $x_{i}\in\mathcal{X}$ represents a conversation between a user and an LLM.
@@
The desired output for our problem is $\bm{s}=\{s_{1},\ldots,s_{k}\}$, consisting of summaries $s_{j}\in\mathcal{S}$. The set $\mathcal{S}$ is potentially infinite (e.g., the set of all strings in our case). The DP guarantee applies only to the release of this output set $\bm{s}$ of summaries.
@@
Evaluating the utility of privacy-preserving summaries presents unique challenges due to the unstructured nature of text data. For utility measurement purposes only, we produce a mapping $\varphi:\bm{x}\to\bm{s}$ so that $\varphi(x_{i})$ is the summary with which the record $x_{i}$ is associated. Importantly, releasing this mapping would violate DP and hence is never returned—it serves solely as an internal mechanism for evaluation.
@@ 2.1 Differential Privacy
All the datasets in this paper are collections of records. A randomized function that maps input datasets to an output space is referred to as a mechanism. For mechanism $\mathcal{M}$ and dataset $\bm{x}$, we use $\mathcal{M}(\bm{x})$ to denote the random variable returned over the output space of $\mathcal{M}$. Two datasets $\bm{x}$ and $\bm{x}^{\prime}$ are said to be adjacent, denoted $\bm{x}\sim\bm{x}^{\prime}$, if, one can be obtained by adding or removing one record from the other; this is referred to as the “add-remove” adjacency. We consider the following notion of $(\varepsilon,\delta)$-differential privacy (DP).
@@
Let $\varepsilon>0$ and $\delta\in[0,1]$. A mechanism $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-differential privacy ($(\varepsilon,\delta)$-DP for short) if for all adjacent datasets $\bm{x}\sim\bm{x}^{\prime}$, and for any (measurable) event $E$ it holds that $\Pr[\mathcal{M}(\bm{x})\in E]\leq e^{\varepsilon}\Pr[\mathcal{M}(\bm{x}^{\prime})\in E]+\delta.$
@@
Clustering is a central primitive in unsupervised machine learning (see, e.g., Ikotun et al., 2023). Abstractly speaking, a (Euclidean) clustering algorithm takes as input a dataset of vectors in $\mathbb{R}^{d}$, and returns a set $C\subseteq\mathbb{R}^{d}$ of $k$ “cluster centers” (for a specified parameter $k$), that are representative of the dataset; formalized for example, via the $k$-means objective (Lloyd, 1982). Clustering algorithms have been well-studied in the literature for several decades, and more recently with DP guarantees as well. We discuss more related work in §E. While our framework is compatible with any DP clustering algorithm, specifically, we use the implementation of DP clustering in the open-source Google DP library [^2] https://github.com/google/differential-privacy/tree/main/learning/clustering , as described in Chang & Kamath (2021); we denote the $(\varepsilon,\delta)$-DP instantiation of any such algorithm producing (up to) $k$ clusters as $\texttt{DP\text{-}K\text{-}Means}_{k,\varepsilon,\delta}$.
@@
In the private histogram release problem over a set $\mathcal{B}$ of bins, the input is a multi-set $\{h_{1},h_{2},\ldots\}$, where each $h_{i}\subseteq\mathcal{B}$ with $|h_{i}|\leq k$. The desired output is a histogram that must be as close to $\sum_{i}h_{i}$ as possible, where we abuse notation to interpret $h_{i}$ as an indicator vector in $\{0,1\}^{\mathcal{B}}$.
@@
In the private partition selection problem over a (potentially infinite) set $\mathcal{B}$, the input is a multi-set $\{h_{1},h_{2},\ldots\}$ where each $h_{i}\subseteq\mathcal{B}$ with $|h_{i}|\leq k$. The desired output is a subset $S\subseteq\mathcal{B}$ such that $S\subseteq\bigcup_{i}h_{i}$ and $S$ is as large as possible. We use the partition selection algorithm introduced by Desfontaines et al. (2022); $\texttt{PPS}_{k,\varepsilon,\delta}$ is the instantiation of this algorithm that is $(\varepsilon,\delta)$-DP. See §A for formal details.
@@ 3 Simple-Clio
We present a non-DP (public) framework for hierarchical text summarization, which is a simplified version of Clio (Tamkin et al., 2024). Our approach transforms a dataset $\bm{x}=(x_{1},\ldots,x_{n})\in\mathcal{X}^{n}$ of conversations into high-level summaries $\bm{s}=(s_{1},\ldots,s_{m})\in\mathcal{S}^{m}$ through three stages. The pseudocode can be found in Algorithm 4 (in §B) with the implementation details as discussed below.
@@
    [Embedding Generation] To use a clustering algorithm, we convert conversations to numerical vectors. A typical way to do this is to use pre-trained embedding models; specifically, we use all-mpnet-base-v2 (Reimers & Gurevych, 2019; 2022). However, following Tamkin et al. (2024), we first map the conversation to a summary text before applying the embedding model; this ensures that conversations that are similar in content but structured differently get mapped to embeddings that are closer. We use an LLM to map each conversation to the summary text, using a prompt as described in §G.1.
@@
    [Clustering] We apply the standard $k$-means clustering (Lloyd, 1982) to group similar conversations based on their embeddings. We select $k$ to achieve an average cluster size of approximately 150 conversations. We assume $\textsc{KMeans}_{k}(\cdot)$ returns the centers of clusters and the assignments (points in each cluster).
@@
    [Summary Generation] For each cluster, we generate a representative summary by sampling both random conversations from the cluster and contrastive conversations near the cluster center. We then use an LLM to produce a summary that captures the main theme of the cluster. The prompt used for this step is provided in §G.2.
This framework provides an approach for organizing and summarizing large collections of conversations, enabling the identification of key topics without relying on manual annotation or predefined categories. However, as mentioned earlier, it does not satisfy any formal privacy guarantee. We next proceed to describing Urania, our DP framework.
4 Urania: A DP Framework for Text Summarization
@@
Input: Input dataset $\bm{x}=(x_{1},\ldots,x_{n})\in\mathcal{X}^{n}$, keyword set $\mathcal{K}$
@@
$\bm{e}\leftarrow\{\mathsf{ExtractEmbeddings}(x_{i}):i\in[n]\}$
@@
    $\bm{K}_{j}\leftarrow\textsf{RandomKeywords}(\mathcal{K},t)$ ;
@@
    $\mathbf{samples}_{j}\leftarrow\textsf{RandomSample}(\bm{C}_{j},m)$ ;
@@
         Append $\textsf{LLMSelectKeywords}(x,\mathcal{K},5)$ to $\mathbf{relevant\_keywords}_{j}$;
@@
    $\bm{K}_{j}\leftarrow\textsf{TopKeywords}(\mathbf{private\_hist}_{j},t)$ ;
@@
$\bm{s}\leftarrow\{\textsf{LLMSummarize}(\bm{K}_{j}):j\in[k]\}$
@@
Algorithm 1 $\textsc{Urania}_{k,t,\varepsilon_{\mathrm{c}},\varepsilon_{\mathrm{hist}},\varepsilon_{\text{size}},\delta_{\mathrm{c}}}$: Differentially Private Text Corpora Summarization
We present Urania, a framework for hierarchical text summarization (with Hierarchy and Simple Visualization presented in §H) that builds on the public framework while providing formal DP guarantees. Unlike Simple-Clio that is non-private, Urania incorporates privacy-preserving mechanisms at critical steps to ensure that the resulting summaries do not leak sensitive information about individual conversations.
Our approach takes a dataset $\bm{x}=(x_{1},\ldots,x_{n})\in\mathcal{X}^{n}$ of conversations, a predefined set $\mathcal{K}$ of keywords, a number $k\in\mathbb{N}$ of clusters, a number of keywords used for summary, and privacy parameters $\varepsilon_{c},\varepsilon_{\mathrm{hist}},\varepsilon_{\mathrm{size}}>0$ and $\delta_{c}\in[0,1]$. The choice of the set $\mathcal{K}$ of keywords is critical to the quality of the generated summaries; we discuss multiple strategies for this in §4.1. Our method produces high-level topical summaries $\bm{s}=(s_{1},\ldots,s_{k})\in\mathcal{S}^{k}$ through the following privacy-preserving pipeline:
@@
    [Embedding Generation] This step is the same as in Simple-Clio.
@@
    [Clustering] Non-private clustering algorithms usually output both cluster centers and cluster assignments. However, cluster assignments are non-private for add-remove adjacency; hence, DP clustering algorithms output only cluster centers.
@@
      First, we generate cluster centers using an $(\varepsilon_{\mathrm{c}},\delta_{\mathrm{c}})$-DP k-means algorithm. Although we use the algorithm described in Chang & Kamath (2021), any DP clustering algorithm can be used here and the same privacy guarantees would follow.
@@
    • Apply DP to the cluster size with privacy parameter $\varepsilon_{\text{size}}$. If the noised size is smaller than a threshold $\tau$, we skip the cluster to avoid generating summaries that will not be meaningful.
@@
    • Use an LLM to extract relevant keywords from the predefined set $\mathcal{K}$ for each sampled conversation. For the purpose of DP, we restrict the number of keywords chosen per conversation to at most 5. See prompts in §G.3.
@@
    [Summary Generation] For each cluster, we generate a high-level summary using only the selected keywords. See prompts in §G.4. We also discuss creating a hierarchy of summaries in §H.
@@
Let $k,t\in\mathbb{N}$, $\varepsilon_{\mathrm{c}},\varepsilon_{\mathrm{hist}},\varepsilon_{\mathrm{size}}>0$, and $\delta_{\mathrm{c}}\in[0,1]$. Then $\textsc{Urania}_{k,t,\varepsilon_{\mathrm{c}},\varepsilon_{\mathrm{hist}},\varepsilon_{\text{size}},\delta_{\mathrm{c}}}$ satisfies $(\varepsilon_{\mathrm{c}}+\varepsilon_{\mathrm{hist}}+\varepsilon_{\mathrm{size}},\delta_{\mathrm{c}})$-DP.
@@
The first part satisfies $(\varepsilon_{\mathrm{c}},\delta_{\mathrm{c}})$-DP since it consists of operations applied per record and subsequent application of the $(\varepsilon_{\mathrm{c}},\delta_{\mathrm{c}})$-DP instantiation of the DP-KMeans algorithm. The second part is $\varepsilon_{\mathrm{size}}$-DP provided that the set of cluster centers is fixed since it again applies per-record operations and an $\varepsilon_{\mathrm{size}}$-DP histogram release algorithm. The third part is $\varepsilon_{\mathrm{hist}}$-DP provided that the set of cluster centers is fixed since it again applies per-record operations and an $\varepsilon_{\mathrm{hist}}$-DP histogram release algorithm, by parallel composition. Altogether, this means that the algorithm producing $\bm{K}_{1}$, …, $\bm{K}_{k}$ is $(\varepsilon_{\mathrm{c}}+\varepsilon_{\mathrm{hist}}+\varepsilon_{\text{size}},\delta_{\mathrm{c}})$-DP due to basic composition.
@@
Remark 4.2.
The process of choosing the set of keywords $\mathcal{K}$ could itself consume some privacy budget, which needs to be accounted for separately. We discuss these methods next.